OpenStudy (anonymous):

Find the point on the graph of z=4x^2−2y^2 at which vector n=<8,8,−1> is normal to the tangent plane. P= ?

OpenStudy (amistre64):

well, lets take the gradient vector

OpenStudy (amistre64):

maybe?

OpenStudy (amistre64):

fx(x,y) = 8x fy(x,y) = -4y

OpenStudy (amistre64):

<8x,-4y> is an equation for the normal to every point (x,y) if i recall correctly

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

8x(x-Px) -4y(y-Py) +0(z-Pz) then... maybe?

OpenStudy (amistre64):

i should prolly include a -1 for the z

OpenStudy (amistre64):

df/dz = -1 so its part of the gradient 8x(x-Px) -4y(y-Py) -1(z-Pz)

OpenStudy (amistre64):

simply put; 8x=8 , and -4y=8

OpenStudy (anonymous):

cross multiply?

OpenStudy (amistre64):

x=2. y=-2

OpenStudy (anonymous):

haha, nm

OpenStudy (amistre64):

put those into your z equation to get the point for z

OpenStudy (amistre64):

z=4(2)^2−2(-2)^2 z = 16 -8 = 8; P(2,-2,8)

OpenStudy (amistre64):

am i right?

OpenStudy (anonymous):

i follow your thought process and agree with it but it says that the first and 3rd coordinates are incorrect

OpenStudy (amistre64):

its possible :)

OpenStudy (anonymous):

haha thats what I said! let me recheck i wrote the correct problem

OpenStudy (amistre64):

we could plug in our normal and try to work it backwards...

OpenStudy (anonymous):

my answer was 2,-2,2

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

8(x-Px)+8(y-Py)-1(z-z0) = 0 look good?

OpenStudy (anonymous):

yes, that looks right

OpenStudy (amistre64):

thats the equation of our tangent plane right?

OpenStudy (amistre64):

z0 = Pz lol mixing memories lol

OpenStudy (anonymous):

yes my book has: z=f(a,b)+fx(a,b)(x-a)+fy(a,b)(y-b)

OpenStudy (amistre64):

2,-2,2 you say was the answer?

OpenStudy (anonymous):

no, thats what I got when i tried to work the problem

OpenStudy (anonymous):

which was also incorrect

OpenStudy (amistre64):

ohhh..... do we have an answer to go by :)

OpenStudy (anonymous):

unfortunetly, no. I do have an example problem similiar to this which I used as model that has an answer

OpenStudy (amistre64):

f(x,y,z) = 4x^2 -2y^2-z = 0 is our original right?

OpenStudy (anonymous):

its z=3x^2-4y^2 n=<3,2,2> and the answer was (-1/4,1/8,1/8)

OpenStudy (anonymous):

yes z=4x^2-2y^2 which is eqivalent to what you said

OpenStudy (amistre64):

and the gradient should give us a parameteric for any normal to any point right?

OpenStudy (amistre64):

<8x,-4y,-1> at any given P(x,y,z)

OpenStudy (amistre64):

z=3x^2-4y^2 n=<3,2,2> <6x,-8y,-1> ( px, py, pz) ------------ <3 , 2, 2> if my thougths are right

OpenStudy (amistre64):

that made no sense lol

OpenStudy (anonymous):

i'm not sure about the gradient, my book doesn't mentioned it. but it says: "vector i normal to the plane: v=<fx(a,b),fy(a,b),-1>

OpenStudy (amistre64):

fx(a,b) is df/dx which is the gradient part... the a,b seems odd to me tho; but they seem to be generic fillins for the inputs that are usually x,y

OpenStudy (amistre64):

your equation has no a or b variables so im assuming it means x,y

OpenStudy (anonymous):

in the example probelm it just converts it fx(x,y)=6x => fx(a,b)=6a

OpenStudy (amistre64):

yeah ....... thats what i was thinking :)

OpenStudy (amistre64):

<6x,-8y,-1> as our 'generic' normal

OpenStudy (amistre64):

<6x,-8y,-1> = <3,2,2> how?

OpenStudy (amistre64):

when x=1/2 and y=-1/4 it appears; but lets see if that makes sense to find z :)

OpenStudy (amistre64):

z = 1/2 when we input those; and we are looking to match: (-1/4,1/8,1/8) (1/2,-1/4,1/2) i wonder if we scale the gradient.... -2<6x,-8y,-1> = <-12x,16y,2> perhaps?

OpenStudy (amistre64):

got it lol

OpenStudy (amistre64):

-12x=3 ; x = -1/4 16y = 2 ; y = 1/8

OpenStudy (amistre64):

8x=8; x=1 .... not 2 lol -4y =8; y=-2 4(1)^2 -2(-2^2) = 4-8 = -4 P(1,2,-4)

OpenStudy (amistre64):

lets see if I can type this without a typo.... P(1,-2,-4) yay!!

OpenStudy (anonymous):

100%!

OpenStudy (anonymous):

you are awesome :)

OpenStudy (amistre64):

thnx :)